OpenStudy (anonymous):

A square piece of cheese covers 8 square inches of a plate. What is the approximate length of one side of the piece of cheese? (Approximate to the nearest hundredth inch.) 2.81 inches 2.82 inches 2.83 inches 2.84 inches

OpenStudy (anonymous):

help please

OpenStudy (arindameducationusc):

yes

OpenStudy (arindameducationusc):

2.82, because

OpenStudy (arindameducationusc):

r^2= 8

OpenStudy (arindameducationusc):

r=length of cheese

OpenStudy (arindameducationusc):

r^2 because (area of square)

OpenStudy (arindameducationusc):

Got it?

OpenStudy (anonymous):

u need tha area of tha cheese area of square=\[b^2\] 8=\[b^2\] 2.82843=b

OpenStudy (triciaal):